OpenStudy (anonymous):

The measure of an angles supplement is 4 times an angles compliment. What's the measure of the supplement?

OpenStudy (radar):

If you know algebra, and you carefully label the information provided for your use by the problem. Let x = the measure of the angle in question. Then we know 180 - x = the supplement of said angle and 90 - x = the compliment of the angle 4(90 - x) = 4 times the compliment. The problem informs you that the supplement is 4 times the compliment. Just set up an equation that states that. See below: 180 - x = 4(90 - x) This states the relationship solve for x. Now the problem is asking for the supplement. After solving for x, the supplement is 180 - x. Solve for the supplement.